OpenStudy (jusaquikie):

how many of these solutions could be used to precipitate Pb2+ ion from .2 M solution of lead nitrate

OpenStudy (jusaquikie):

i think they would break down as follows Lead Nitrate Pb2+ 2(NO3) KCl = K+ Cl- KOH = K+ (OH)- NaCH3COO = guessing Na+ CH3- COO CaBr2 = Ca^2+ Br-

OpenStudy (jusaquikie):

The answer is 3 but i don't know how to figure this out. I would have guessed 1 because CaBr2 is the only 2Br- to strip away the2 (NO3)

OpenStudy (aaronq):

NaCH3COO would dissociate in to Na+ and CH3COO- (which is the conjugate base of acetic acid) you're finding out which solution would precipitate Pb^2+ out so look at solubility tables, i'm pretty sure Pb(OH)2 is insoluble maybe PbBr2

OpenStudy (jusaquikie):

ok on the soulubility chart provided exceptions are Br- Cl- I- and (SO4)2-

OpenStudy (jusaquikie):

does that mean PbBr would be insouluble so that would be a precpitate?

OpenStudy (aaronq):

PbBr2 ..since Pb is 2+ and Cl is 1- .. balance out charges the chart i have says PbBr2, PbCl2 and Pb(OH)2 are insoluble

OpenStudy (aaronq):

so yeah, only the third one would not cause a precipitate

OpenStudy (jusaquikie):

and (OH)- is insoulible with no exceptions for Pb so that would make 3 this makes sense now, i was trying to have Pb2+ left over after froming something witht the rest of the ions.

OpenStudy (jusaquikie):

so all but the NaCH3COO

OpenStudy (aaronq):

yeah, thats all you can do to show they will precipitate. if you were given more information (like the concentration of the other solutions) you could've shown they precipitate using their Ksp's

OpenStudy (jusaquikie):

Thank you again for your time i will be giving out medals all night lol i have my chem final in the morning =p Thank you

OpenStudy (aaronq):

no problem man, i have those next week also. good luck with your final !